OpenStudy (popstar30977):

660 ft/min__________________ft/s round to the nearest tenth if necessary Please helpT_T

OpenStudy (anonymous):

Well you know that there are 60 seconds in a minute so all you need to do is multiply 660 times 60 and you will your answer I believe

OpenStudy (anonymous):

660×60=39,600

OpenStudy (anonymous):

@satellite73 do you agree?

OpenStudy (nevermind_justschool):

Please stop spamming the chat since lovely person is trying to help you :)) Best of luck :)

OpenStudy (anonymous):

660ft/min=660/60 ft/s

OpenStudy (anonymous):

Thank you @nevermind_justschool

OpenStudy (anonymous):

@surjithayer why would you divide when there are going to be more seconds then there are minutes?

OpenStudy (anonymous):

if a moving body covers distance 660 ft in one min.,how much it moves in one second (more or less)

OpenStudy (popstar30977):

25mi/h=____________ft/s

OpenStudy (popstar30977):

^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

Oh ok so I guess I was wrong then. 660/60=11 sorry @popstar30977

OpenStudy (popstar30977):

it is ok @kimbllse777

OpenStudy (anonymous):

\[25mile/hr=\frac{ 25 \times 1760 \times 3 }{ 60 \times 60 }ft/s\]

OpenStudy (anonymous):

Did you solve the equation that surji gave you @Popstar30977?

OpenStudy (popstar30977):

Yes thank you @kimbllse777